OpenStudy (aayushi.somani):

At what height above the earth's surface, the value of g is same as in a mine 80 km deep?

OpenStudy (aayushi.somani):

@sarah786

OpenStudy (anonymous):

at height variation of graivty is g '=g(1-[(2h)/R]) and in depth g '=g(1-[d/R]) so (1-[(2h)/R])=(1-[d/R]) where R= radius od earth and d=80km(given) you have to find 'h '. calculate and solve Or it will cme like this... g(1-2h/R) = g(1-d/R) => 1-2h/R = 1- d/R =>2h/R = d/R => 2h = d => 2h = 80 => h = 40 m rest things are given above..... Source: http://www.meritnation.com/

OpenStudy (aayushi.somani):

thankq @sarah786

OpenStudy (anonymous):

you're welcome but I just google'd it and I found all that , Did ya ???

OpenStudy (aayushi.somani):

noo i just popsted it directly..! :P

OpenStudy (anonymous):

lol ok then :P